An important radioisotope for industrial and medical applications is 60Co with a half life of 5.2 years and a molar mass of 60 grams/mole. What minimum initial mass of 60Co is required for a sealed cource of 60Co to have an activity of 10 Ci after 30 months where 1Ci=3.7x1010 Bq. 


1
Expert's answer
2021-06-11T11:21:02-0400

Convert from Ci to Bq: 10 Ci is 3.7·1011 Bq.


Determine decay constant:


"\lambda=\frac{\ln2}{t}=0.0111\text{ month}^{-1}."

Find initial activity:


"A=A_0e^{-\lambda t},\\\space\\
A_0=\frac{A}{e^{-\lambda t}}=5.16·10^{11}\text{ Bq}."

Activity can be expressed as a product of decay constant and number of particles:


"A=\lambda N."

Find initial N:


"N_0=\frac{A_0}{\lambda}=1.22·10^{22}."

Find the amount of substance:


"n_0=\frac{N_0}{N_A}=0.0203\text{ mol}."

Find the mass:


"m_0=n_o\Mu=1.22\text{ g}."